# Safe Reinforcement Learning In Black-Box Environments Via Adaptive Shielding

### v2

@@ Appendix A Task Details
In the Safety Gymnasium [Ji et al., 2023b]  [^2] More details here: https://github.com/PKU-Alignment/safety-gymnasium test-suite, a robot with Lidar sensors has to navigate through environments with obstacles to complete a given task. The test suite comes with a multitude of robots (e.g. Point, Ant, Car) and a set of tasks (e.g. Goal, Circle, Button) that can be evaluated. In our experiments, found in Section 6, we chose to use:
@@
  • Semi-random Goal: A standard goal environment, where the agent aims to reach the goal at the end of the episode whilst navigating through six obstacles. The six obstacles have a static spawn, the agent and the goal have randomised positions every episode. We placed the six obstacles to form a large wall, where the agent can fit through to reach the goal but with an increased risk of crashing. In this instance, we want to determine if ADVICE and other safe RL exploration methods will learn to avoid the wall or risk going through it. Deployment trajectories in Figure 3, show the learnt trajectories of ADVICE and other methods.
@@
  • Constrained Randomised Goal: This environment is similar to the random goal environment, but instead the task is set up as a constrained MDP. Instead of obstacles that terminate the episode, the task includes hazards that give the agent $-0.2$ cost per step when inside them. The agent cannot terminate in this task and instead has to minimise cost whilst maximizing rewards.
In all environments for all the tasks above, the agent uses psuedo Lidar to perceive objects in the environment. Each type of object (e.g. goal, obstacles) in the environment has its own separate Lidar observation, where a Lidar vector has 16 bins. All vectors are flattened into one observational vector and then given to the agent as the current state. For example, in the semi-random goal environment, there is a goal and a set of obstacles. Here the observational space is $\left[0,1\right]^{32}$. All lidar vectors have a max distance of $3$ meters. In both goal environments, an episode has a maximum timestep of $1000$. In the circle environment, the maximum timestep is $500$. A goal reach in both goal environments is defined as reaching the goal before the episode truncates, in the circle environment it is defined as being within the circle when the episode truncates. A crash is defined as the agent colliding with an obstacle, when this occurs, the episode is terminated shortly. Anything else is considered as the episode timing out.
@@
Figure 5: Different views of the Car robot in the Safety Gymnasium test suite [Ji et al., 2023a].
Each task has a separate reward function for the agent to maximise. Whenever the agent comes into contact with an obstacle, a constraint cost of $-1$ is given (the exception being the Constrained Randomised Goal task). The reward functions for each task are:
  • Semi-random, Randomised Goal, Randomised Button, & Constrained Randomised Goal: $R_{t}=(D_{last}-D_{now})\beta$, where $D_{last}$ is the distance between the agent and the goal in timestep $t_{-1}$, $D_{now}$ is the distance between the agent and the goal in timestep $t$, and $\beta$ is the discount factor. Simply, the agent moving towards the goal, in terms of Euclidean distance, gains a positive reward. The agent moving away from the goal gains a negative reward. Reaching the goal gains a static reward of $+1$.
It should be emphasized that the agent or methods used in the evaluation have no prior knowledge of the task/environment/safety concern. Thus, we can define this environment and all tasks within as black box.
(a) Semi-random
Goal
(b) Randomised
Goal
(c) Randomised
Button
(d) Randomised
Circle
(e) Constrained
Randomised Goal
Figure 6: Example navigation tasks with varying complexity levels for evaluating ADVICE. The purple blocks are the obstacles, the green circles are the goals, orange spheres are the buttons, and the red vehicle is the agent.
@@ B.1 Theorem 1
Theorem 1. The probability of ADVICE misclassifying a feature is bounded by $exp(-\gamma/2\sigma^{2})$, where $\gamma$ is the contrastive separation margin and $\sigma^{2}$ is the variance of the assumed Gaussian noise in the latent space.
@@
$\gamma=\min_{f_{s}\in S,f_{u}\in U}\|\textsc{Encode}(f_{s})-\textsc{Encode}(f_{u})\|_{2}$ (6)
@@
Let $d_{s}=min_{f_{s}\in S}\|\textsc{Encode}(f),\textsc{Encode}(f_{s})\|_{2}$ and $d_{u}=min_{f_{u}\in U}\|\textsc{Encode}(f),\textsc{Encode}(f_{s})\|_{2}$ define the Euclidean distance to the nearest safe and unsafe feature, respectively, given a new feature $f$.
An incorrect decision (misclassification) occurs when $f\in S$ and $d_{u}<d_{s}$ or $f\in U$ and $d_{u}>d_{s}$. The contrastive separation margin $\gamma$ ensures that, in a noise-free case $\|\textsc{Encode}(f_{u}),\textsc{Encode}(f_{s})\|_{2}\geq\gamma\;\forall f_{s}\in S\;\forall s_{u}\in U$. In RL environments such as those targeted by ADVICE where noise exists, the distances $d_{s}$ and $d_{u}$ are perturbed by $\epsilon_{s},\epsilon_{u}\sim N(0,\sigma^{2})$. Therefore:
@@
where $\epsilon=\epsilon_{u}-\epsilon_{s}\sim N(0,\sigma^{2})$ and is independent and Gaussian. The probability of $d_{u}<d_{s}$ (misclassification) when $f\in S$ is given by:
@@
ADVICE uses K-nearest neighbours to classify an unseen feature $f$. If $\gamma$ is large relative to $\sigma$, the probability of misclassifying an unseen feature decreases exponentially. So, we can define the expected number of misclassified features to be:
$E[misclassified\;features]\leq N\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)$ (11)
The noise $\sigma$ in the latent space can come from: noisy data, imperfect model training, randomness in batch sampling, etc. ∎
@@ B.2 Proof of Theorem 2
Let $\gamma_{m}$ express the effective achieved separation margin between sets $S$ and $U$, where $\gamma_{m}\leq\gamma$. Equality only holds under ideal conditions, such as perfectly diverse data, perfect model training, no data noise, etc. The diversity of the feature set $F_{E}$ collected before episode $E$ can be quantified using entropy:
@@
where $p(f)$ is the probability distribution of features $f\in F_{E}$. Higher entropy corresponds to a broader set of features, ensuring greater diversity. Greater diversity results in more representative embeddings, given good model training, allowing the contrastive loss function to achieve better separation and clusterings of sets $S$ and $U$. The effective separation margin $\gamma_{m}$ depends on $H(F_{E})$. As diversity increases, the embeddings for $S$ and $U$ become more separable, thereby $\gamma_{m}\;\propto\;k\cdot\sqrt{H(F_{E})}$ where $k>0$ is a proportionality factor that links the entropy $H(F_{E})$ of the feature set to the effective separation margin $\gamma_{m}$. It encapsulates the influence of latent space geometry, scaling properties, and model-specific parameters. While $k$ may vary depending on the training process and feature distribution, it is assumed to be stable for a given setup. Empirically, $k$ can be estimated by observing the relationship between $\gamma_{m}$ and $H(F_{E})$ across diverse datasets or configurations. From Theorem 1, with the substitution of $\gamma_{m}$ for $\gamma$, the probability of misclassifying a feature is bounded by:
@@ B.3 Proof of Theorem 3
Theorem 3. The expected regret of ADVICE with a fixed $K$ is bounded by as $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$.
@@
The total regret of ADVICE can be decomposed into two individual components: the baseline regret from the reinforcement learning process, and regret introduced from the shielding mechanism with a fixed $K$ accounting for misclassification. The baseline regret arises from the agent’s standard exploration-exploitation trade-off in reinforcement learning. This term is defined as:
$=O\left(\frac{1}{\sqrt{E_{max}}}\right)$ (14)
which reflects the improvement in the agent’s policy quality over $E_{max}$ episodes. To construct the regret term for regret introduced from the shielding mechanism with a fixed $K$ accounting for misclassification, we can use the probabilistic bound of misclassification from Theorem 1, $exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)$. Let $\Delta Q=Q^{*}(s_{t},a_{t})-Q^{*}(s_{t},a^{{}^{\prime}}_{t})$ denote the average regret per timestep caused by taking the misclassified action $a^{{}^{\prime}}_{t}$ instead of the optimal action $a_{t}$. The shield also restricts the agent’s action choices into a feasible set $\mathcal{F_{K}}$ determined by $K$ where: a higher $K$ (more cautious shield) reduces $\mathcal{F_{K}}$, whilst a lower $K$ (more relaxed shield) increases $\mathcal{F_{K}}$. Thus the size of $\mathcal{F_{K}}$ scales inversely with $\sqrt{K}$. We can define our regret for the shielding term as:
$=O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$ (15)
@@ B.4 Proof of Theorem 4
Theorem 4. The expected long-term regret of ADVICE with an adaptive $K$ is lower than ADVICE with a fixed $K$ and is bounded by $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)+O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$.
@@
$=O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$ (16)
@@
$=O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)$ (18)
@@
While adaptive $K$ introduces early added regret but diminishes as $K_{e}\rightarrow K^{*}$. This ensures that the long-term regret is dominated by the optimal $K^{*}$, unlike a fixed $K$ which does not adapt to the environment. Let $\epsilon$ represent the difference between $K$ and $K^{*}$. For a fixed $K$, $\epsilon$ is constant leading to regret terms proportional to $|\epsilon|$. For adaptive $K$, $|\epsilon_{e}|$ decreases over episodes due to convergence $|\epsilon_{e}|\propto\frac{1}{\sqrt{e}}$. This makes the additional regret introduced by adaptive $K$ approach 0 as $E_{max}\rightarrow\infty$. ∎
Appendix C Full Training Results
In Section 6.1, we show the main results for all methods in a set of tasks. For fair comparison, we show results from episode $1000$ and standardise all metrics to zero. Below, in Figure 7, we show the unstandardised results for the same experiments.
@@ Appendix C Full Training Results
Figure 7: Average episodic reward, cumulative safety violations, cumulative goal reaches of examined methods (DDPG, DDPG-Lag, RCPO, Tabular shield, Conservative Safety Critic, ADVICE) and example latent space visualisation for the semi-random goal (a), randomised goal (b), randomised button (c), and randomised circle (d) environments.
@@
In all experiments, the Tabular Shield method performs approximately the same as the standard DDPG agent. To show why this behaviour occurs, we plot the average shield activations for ADVICE and the Tabular Shield in Figure 9. From these results, it is evident that the Tabular Shield does not once activate during training across all tasks. This is due to the high dimensionality of the environments evaluated. Even though the features stored are discretised to $1$ decimal place, the agent has to observe the exact same values across all $\approx 32$ dimensions plus the actions for the shield to activate. Our experiments show that this method fails in these types of environments. A trend that can be noticed with ADVICE is when the shield is first activated, the amount of interventions starts relatively high. As training progresses this number reduces, which shows that the agent learns to adapt to the shield’s understanding of safety.
Figure 10: The rate of safety violations for all methods in the semi-random goal, random goal, random button, and circle environments respectively.
@@
While constrained MDP environments are less common than unconstrained ones, we evaluated ADVICE in the Constrained Randomized Goal environment in Figure 8. ADVICE consistently achieved one of the lowest average episodic costs, performing on par with cost-based methods, which showed significant oscillations at times. These fluctuations highlight the difficulty these methods face in balancing cost reduction with reward maximization [Liu et al., 2022]. Notably, such oscillations were absent in sparse-cost environments like the unconstrained settings in Section 6.1.

### v3

@@ Appendix A Task Details
In the Safety Gymnasium [25]  [^2] More details here: https://github.com/PKU-Alignment/safety-gymnasium test-suite, a robot with Lidar sensors has to navigate through environments with obstacles to complete a given task. The test suite comes with a multitude of robots (e.g. Point, Ant, Car) and a set of tasks (e.g. Goal, Circle, Button) that can be evaluated. In our experiments, found in Section 6, we chose to use:
@@
  • Semi-random Goal: A standard goal environment, where the agent aims to reach the goal at the end of the episode whilst navigating through six obstacles. The six obstacles have a static spawn, the agent and the goal have randomised positions every episode. We placed the six obstacles to form a large wall, where the agent can fit through to reach the goal but with an increased risk of crashing. In this instance, we want to determine if ADVICE and other safe RL exploration approaches will learn to avoid the wall or risk going through it. Deployment trajectories in Figure 3, show the learnt trajectories of ADVICE and other approaches.
@@
  • Constrained Randomised Goal: This environment is similar to the random goal environment, but instead the task is set up as a constrained MDP. Instead of obstacles that terminate the episode, the task includes hazards that give the agent $0.2$ cost per step when inside them. The agent cannot terminate in this task and instead has to minimise cost whilst maximizing rewards.
In all environments for all the tasks above, the agent uses psuedo Lidar to perceive objects in the environment. Each type of object (e.g. goal, obstacles) in the environment has its own separate Lidar observation, where a Lidar vector has 16 bins. All vectors are flattened into one observational vector and then given to the agent as the current state. For example, in the semi-random goal environment, there is a goal and a set of obstacles. Here the observational space is $\left[0,1\right]^{32}$. All lidar vectors have a max distance of $3$ meters. In both goal environments, an episode has a maximum timestep of $1000$. In the circle environment, the maximum timestep is $500$. A goal reach in both goal environments is defined as reaching the goal before the episode truncates, in the circle environment it is defined as being within the circle when the episode truncates. A crash is defined as the agent colliding with an obstacle, when this occurs, the episode is terminated shortly. Anything else is considered as the episode timing out. For the cost-based approaches in the unconstrained environments, the approaches receive a cost of $0.2$ when entering a non-accepting terminal state. This was chosen via a mixture of manual tuning, the recommended values from Safety Gymnasium, and comparative approaches.
Each task has a separate reward function for the agent to maximise. Whenever the agent comes into contact with an obstacle, a constraint cost of 1-1 is given (the exception being the Constrained Randomised Goal task). The reward functions for each task are:
  • Semi-random, Randomised Goal, Randomised Button, & Constrained Randomised Goal: $R_{t}=(D_{last}-D_{now})\beta$, where $D_{last}$ is the distance between the agent and the goal in timestep $t_{-1}$, $D_{now}$ is the distance between the agent and the goal in timestep $t$, and $\beta$ is the discount factor. Simply, the agent moving towards the goal, in terms of Euclidean distance, gains a positive reward. The agent moving away from the goal gains a negative reward. Reaching the goal gains a static reward of +1+1.
It should be emphasized that the agent or approaches used in the evaluation have no prior knowledge of the task/environment/safety concern. Thus, we can define this environment and all tasks within as black box.
@@
Figure 7: Different views of the Car robot in the Safety Gymnasium test suite [24].
(a) Semi-random Goal
(b) Randomised Goal
(c) Randomised Button
(d) Randomised Circle
(e) Constrained Randomised Goal
Figure 8: Example navigation tasks with varying complexity levels for evaluating ADVICE. The purple blocks are the obstacles, the green circles are the goals, orange spheres are the buttons, and the red vehicle is the RL agent.
@@ B.1 Theorem 1
Theorem 1. The probability of ADVICE misclassifying a feature is bounded by $exp(-\gamma^{2}/2\sigma^{2})$, where $\gamma$ is the contrastive separation margin and $\sigma^{2}$ is the variance of the assumed Gaussian noise in the latent space.
@@
$\gamma=\min_{f_{s}\in F_{safe},f_{u}\in F_{unsafe}}\|\textsc{Encode}(f_{s})-\textsc{Encode}(f_{u})\|_{2}$ (6)
@@
Let $d_{s}=min_{f_{s}\in F_{safe}}\|\textsc{Encode}(f),\textsc{Encode}(f_{s})\|_{2}$ and $d_{u}=min_{f_{u}\in F_{unsafe}}\|\textsc{Encode}(f),\textsc{Encode}(f_{s})\|_{2}$ define the Euclidean distance to the nearest safe and unsafe feature, respectively, given a new feature $f$.
An incorrect decision (misclassification) occurs when $f\in F_{safe}$ and $d_{u}<d_{s}$ or $f\in F_{unsafe}$ and $d_{u}>d_{s}$. The contrastive separation margin $\gamma$ ensures that, in a noise-free case $\|\textsc{Encode}(f_{u}),\textsc{Encode}(f_{s})\|_{2}\geq\gamma\;\forall f_{s}\in F_{safe}\;\forall s_{u}\in F_{unsafe}$. In RL environments such as those targeted by ADVICE where noise exists, the distances $d_{s}$ and $d_{u}$ are perturbed by $\epsilon_{s},\epsilon_{u}\sim N(0,\sigma^{2})$. Therefore:
@@
where $\epsilon=\epsilon_{u}-\epsilon_{s}\sim N(0,\sigma^{2})$ and is independent and Gaussian. The noise $\sigma$ in the latent space can come from: noisy data, imperfect model training, randomness in batch sampling, etc. The probability of $d_{u}<d_{s}$ (misclassification) when $f\in S$ is given by:
@@
ADVICE uses K-nearest neighbours to classify an unseen feature $f$. If $\gamma$ is large relative to $\sigma$, the probability of misclassifying an unseen feature decreases exponentially. So, we can define the expected probability of misclassifying a feature to be:
$E[\textrm{feature misclassification}]\leq exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)$ (11)
@@ B.2 Proof of Theorem 2
Let $\gamma_{m}$ express the effective achieved separation margin between sets $F_{safe}$ and $F_{unsafe}$, where $\gamma_{m}\leq\gamma$. Equality only holds under ideal conditions, such as perfectly diverse data, perfect model training, no data noise, etc. The diversity of the feature set $F_{E}$ collected before episode $E$ can be quantified using entropy:
@@
where $p(f)$ is the probability distribution of features $f\in F_{E}$. Higher entropy corresponds to a broader set of features, ensuring greater diversity. Greater diversity results in more representative embeddings, given good model training, allowing the contrastive loss function to achieve better separation and clusterings of sets $F_{safe}$ and $F_{unsafe}$. The effective separation margin $\gamma_{m}$ depends on $H(F_{E})$. As diversity increases, the embeddings for $F_{safe}$ and $F_{unsafe}$ become more separable, thereby $\gamma_{m}\;\propto\;k\cdot\sqrt{H(F_{E})}$ where $k>0$ is a proportionality factor that links the entropy $H(F_{E})$ of the feature set to the effective separation margin $\gamma_{m}$. It encapsulates the influence of latent space geometry, scaling properties, and model-specific parameters. While $k$ may vary depending on the training process and feature distribution, it is assumed to be stable for a given setup. Empirically, $k$ can be estimated by observing the relationship between $\gamma_{m}$ and $H(F_{E})$ across diverse datasets or configurations. From Theorem 1, with the substitution of $\gamma_{m}$ for $\gamma$, the probability of misclassifying a feature is bounded by:
@@ B.3 Proof of Theorem 3
Theorem 3. The expected regret of ADVICE with a fixed $K$ is upper bounded by a function of order $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$.
@@
The total regret of ADVICE can be decomposed into two individual components: the baseline regret from the reinforcement learning process, and regret introduced from the shielding mechanism with a fixed $K$ accounting for misclassification. The baseline regret arises from the agent’s standard exploration-exploitation trade-off in reinforcement learning [10]. This term is defined as:
$O\left(\frac{1}{\sqrt{E_{max}}}\right)$ (14)
which reflects the improvement in the agent’s policy quality over $E_{max}$ episodes. To construct the regret term for regret introduced from the shielding mechanism with a fixed $K$ accounting for misclassification, we can use the probabilistic bound of misclassification from Theorem 1, $exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)$. Let $\Delta Q=\sum_{t=1}^{T}\frac{Q^{*}(s_{t},a_{t})-Q^{*}(s_{t},a^{{}^{\prime}}_{t})}{T}$ denote the average regret per timestep caused by taking the misclassified action $a^{{}^{\prime}}_{t}$ instead of the optimal action $a_{t}$. The shield also restricts the agent’s action choices into a feasible set $\mathcal{F_{K}}$ determined by $K$ where: a higher $K$ (more cautious shield) reduces $\mathcal{F_{K}}$, whilst a lower $K$ (more relaxed shield) increases $\mathcal{F_{K}}$. Thus, the size of $\mathcal{F_{K}}$ scales inversely with $\sqrt{K}$. We can define our regret for the shielding term as:
$O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K}}\right)$ (15)
@@ B.4 Proof of Theorem 4
Theorem 4. The expected long-term regret of ADVICE with an adaptive $K$ is lower than ADVICE with a fixed $K$ and is upper bounded by a function of order $O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)+O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$.
@@
$O\left(\frac{1}{E_{max}}\sum^{E_{max}}_{e=1}|K_{e}-K^{*}|\cdot\Delta Q_{K_{e}}\right)$ (16)
@@
$O\left(\frac{1}{\sqrt{E_{max}}}\right)+O\left(E_{max}\cdot\Delta Q\cdot exp\left(-\frac{\gamma^{2}}{2\sigma^{2}}\right)\cdot\frac{1}{\sqrt{K^{*}}}\right)$ (18)
@@
While adaptive $K$ introduces early added regret, but diminishes as $K_{e}\rightarrow K^{*}$. This ensures that the long-term regret is dominated by the optimal $K^{*}$, unlike a fixed $K$, which does not adapt to the environment. Let $\epsilon$ represent the difference between $K$ and $K^{*}$. For a fixed $K$, $\epsilon$ is constant leading to regret terms proportional to $|\epsilon|$. For adaptive $K$, $|\epsilon_{e}|$ decreases over episodes due to convergence $|\epsilon_{e}|\propto\frac{1}{\sqrt{e}}$. This makes the additional regret introduced by the adaptive $K$ approach 0 as $E_{max}\rightarrow\infty$. ∎
@@
Figure 9: Average episodic reward, cumulative safety violations, cumulative goal reaches of examined approaches (DDPG, DDPG-Lag, RCPO, Tabular shield, Conservative Safety Critic, ADVICE) and example latent space visualisation for the semi-random goal (a), randomised goal (b), randomised button (c), and randomised circle (d) environments.
Appendix C Full Training Results
In Section 6.1, we show the main results for all approaches in a set of tasks. For fair comparison, we show results from episode $1000$ and standardise all metrics to zero. Below, in Figure 9, we show the unstandardised results for the same experiments.
@@ Appendix C Full Training Results
In all experiments, the Tabular Shield approach performs approximately the same as the standard DDPG agent. To show why this behaviour occurs, we plot the average shield activations for ADVICE and the Tabular Shield in Figure 11. From these results, it is evident that the Tabular Shield does not activate once during training across all tasks. This is due to the high dimensionality of the environments evaluated. Even though the features stored are discretised to $1$ decimal place, the agent has to observe the exact same values across all $\approx 32$ dimensions plus the actions for the shield to activate. Our experiments show that this approach fails in these types of environments. A trend that can be noticed with ADVICE is when the shield is first activated, the amount of interventions starts relatively high. As training progresses, this number reduces, which shows that the agent learns to adapt to the shield’s understanding of safety.
Figure 12: The rate of safety violations for all approaches in the semi-random goal, random goal, random button, and circle environments, respectively.
@@
While constrained MDP environments are less common than unconstrained ones, we evaluated ADVICE in the Constrained Randomized Goal environment in Figure 10. ADVICE consistently achieved one of the lowest average episodic costs, performing on par with cost-based approaches, which showed significant oscillations at times. These fluctuations highlight the difficulty these approaches face in balancing cost reduction with reward maximization [30]. Notably, such oscillations were absent in sparse-cost environments like the unconstrained settings in Section 6.1.